Can you please show me step by step how to solve the following;
Solve by completing the square
x^2+8x+13=0

Answer by scott8148(6628)

a quadratic (ax^2+bx+c) is a "perfect" square if (b^2)/4a=c

subtract 13 to clear the c value __ x^2+8x=-13

add (b^2)/4a __ x^2+8x+16=3 __ (x+4)^2=3

take sqrt __ x+4=ħsqrt(3)

subtract 4 __ x=-4ħsqrt(3)